# Multi-Modal Optimization with k-Cluster Big Bang-Big Crunch Algorithm

### v1

@@ Abstract
Multi-modal optimization is often encountered in engineering problems, especially when different and alternative solutions are sought. Evolutionary algorithms can efficiently tackle multi-modal optimization thanks to their features such as the concept of population, exploration/exploitation, and being suitable for parallel computation.
This paper introduces a multi-modal optimization version of the Big Bang-Big Crunch algorithm based on clustering, namely, k-BBBC. This algorithm guarantees a complete convergence of the entire population, retrieving on average the $99\%$ of local optima for a specific problem. Additionally, we introduce two post-processing methods to (i) identify the local optima in a set of retrieved solutions (i.e., a population), and (ii) quantify the number of correctly retrieved optima against the expected ones (i.e., success rate).
@@
Our results show that k-BBBC performs well even with problems having a large number of optima (tested on $379$ optima) and high dimensionality (tested on $32$ decision variables). When compared to other multi-modal optimization methods, it outperforms them in terms of accuracy (in both search and objective space) and success rate (number of correctly retrieved optima) – especially when elitism is applied. Lastly, we validated our proposed post-processing methods by comparing their success rate to the actual one. Results suggest that these methods can be used to evaluate the performance of a multi-modal optimization algorithm by correctly identifying optima and providing an indication of success – without the need to know where the optima are located in the search space.
Index Terms:
Multi-Modal Optimization, Evolutionary Algorithms, Clustering, Big Bang-Big Crunch.
I Introduction
@@
Mathematical optimization is the search for the best element within a set of alternatives based on specific criteria, with applications that range in many different areas such as science, economics, and engineering [1, 2, 3, 4, 5]. The process usually consists of exploring the search space of a given objective function, which could be limited by well-defined constraints, to retrieve a single global optimal solution that either minimizes or maximizes the objective – namely, global optimization. When the objective function features more than one suitable solution (i.e., optima or modes), the process is known as multi-modal optimization.
Many practical optimization problems have more than one optimal solution – whether that is multiple optima with the same objective value or a single global optimum surrounded by many local optima. In such cases, retrieving a single optimum might not be sufficient to tackle the problem due to different reasons:
  • a global optimum might be defined by parameters that are unrealistic or excessively expensive [6];
  • a solution can be optimal based on the availability of essential resources or the fulfillment of specific criteria, but if these conditions change over time, users need to adapt to different solutions [7];
  • knowing various ideal solutions within the search space can offer a valuable understanding of the characteristics of the problem and its solutions [8];
  • robust solutions are preferred over global ones because if the optimal solution lies on a peak of the function, any small change in its value would cause instability to the system, whereas a sub-optimal solution lying in a wider area or a plateau is less susceptible to change [9]; and
  • reliable solutions are preferred over global ones, as uncertainty in the search space might lead to infeasible optimal solutions (i.e., in violation of constraints) [10].
Under these circumstances, a local optimum offering decent performance at a reasonable cost might be more desirable than a costly global optimum providing a marginal better performance [11].
@@
Evolutionary Computation offers a set of optimization algorithms, namely Evolutionary Algorithms (EAs), which can easily tackle complex optimization problems such as multi-modal ones. Thanks to their stochastic exploration of space, along with the inherent parallel computation offered by their population nature, they can easily locate different good solutions at the same time [12, 13]. Many multi-modal evolutionary algorithms (MMEAs) have been proposed in the literature, featuring diverse strategies such as: (i) implicitly promote diversity by spatial segregation [14, 15] or spatial distribution [16, 17, 18, 19, 20, 21] of a single population; (ii) partitioning the population [22] and imposing mating restrictions between different populations [23, 24] or depending on the individuals’ life cycle [25]; (iii) with elitist methods [26] or methods that conserve the genotype of individuals [27]; (iv) enforcing diversity with fitness sharing [28, 29, 30], or clearing [31], crowding [32, 33, 34, 35, 36, 37], clustering [38]; and (v) employing multi-objective optimization techniques [7]. However, most of these methods suffer from difficulties such as tuning niching parameters, poor scalability with high dimensionality, extra computational overhead, not being able to fully explore the search space, relying on gradient information, premature convergence, and failure to find the right balance between exploration (finding new solutions in the search space) and exploitation (making use of the features of the good solutions discovered throughout the run) [39, 40].
@@
In this work, we propose an innovative MMEA that extends the Big Bang-Big Crunch algorithm [41] to solve multi-modal optimization problems. Our algorithm, namely the k-Cluster Big Bang-Big Crunch algorithm (k-BBBC), is a cluster-based method that relies on uniform exploration of the search space, as well as on exploitation of its best solutions. One of the strengths of BBBC is to dynamically turn exploration into exploitation during a single run, guaranteeing convergence to an optimal solution; we used this property allowing k-BBBC to correctly and fully converge to every local optimum of a problem – note that the entire population will converge. However, the amount of initial exploration must be sufficient to find all the optima [42, 43]. To ensure a proper amount of exploration, we mathematically defined the relationship between the population size and the number of local optima to be retrieved (when known), which allowed k-BBBC to find all optima for every problem we tested – with some minor exceptions. When compared to other MMEAs, we found that k-BBBC outperforms them in terms of accuracy and number of retrieved solutions. Note that k-BBBC works for both minimization and maximization problems, and features both non-elitist and elitist versions.
@@
However, returning a population of converged solutions without identifying which ones are the optima is not efficient (i.e., we cannot simply sort solutions based on their fitness). We certainly do not want to manually compare solutions against each other: this quickly becomes unfeasible when the population size and the number of expected optima increases. Therefore, we propose efficient and automated methods for post-processing the output of an MMEA – theoretically, every MMEA provided that its population fully converged. The methods include (i) an identification procedure to detect the local optima in a population of converged solutions, and (ii) a quantification procedure to determine how many local optima have been missed (when this information is available). All these methods are based on clustering.
The rest of the paper is organized as follows: Sec. II describes k-BBBC; Sec. III describes post-processing procedures for precise localization of optimal solutions within a wide population of individuals; Sec. IV illustrates a set of experiments to compare k-BBBC with other MMEAs; Sec. V discusses the advantages and disadvantages of k-BBBC based on the experimental results; and finally, Sec. VI summarizes the work and presents possible future ideas.
II Optimization Method
@@
This section will present our proposed multi-modal optimization method, starting from its original version for global optimization to a full walk-through of our proposed method and its behavior. We will be using the function defined in (1), namely the Key function, to provide visual reference while describing the algorithm. This function features a single decision variable $x$ and a changeable number of optima depending on the parameter $m$ – featuring only one global minimum. Lastly, Table I reports the list of the algorithm’s parameters along with the symbol we will use to refer to them in the following sections.
@@
TABLE I: Algorithm’s Parameters
@@
$g$ max number of iterations/generation of the algorithm
$m$ number of local optima of the problem (if known)
$k$ number of clusters dividing the population
@@
$d$ problem dimensionality (i.e., number of decision variables)
II-A Big Bang-Big Crunch Algorithm for Global Optimization
@@
BBBC features different crunch operators [44], which determine how the center of mass $c$ is calculated. It can be done by weighting the individuals with corresponding fitness evaluations as in (2), or by choosing the most fitting individual as in (3).
$c=\frac{\sum_{i=1}^{n}{\frac{1}{f(\textbf{x}_{i})}\textbf{x}_{i}}}{\sum_{i=1}^{n}{\frac{1}{f(\textbf{x}_{i})}}},\quad\textbf{x}_{i}\in P$ (2)
$c=\begin{dcases}\textbf{x}_{j}:f(\textbf{x}_{j})=\max_{i=1}^{n}{f(\textbf{x}_{i})}\;\text{if maximization}\\ \textbf{x}_{j}:f(\textbf{x}_{j})=\min_{i=1}^{n}{f(\textbf{x}_{i})}\;\text{if minimization}\end{dcases},\;\textbf{x}_{i}\in P$ (3)
The state-of-the-art reports only one type of bang operator [41], which generates new individuals $\textbf{x}^{new}$ for the next population as expressed in (4). The center of mass $c$ is noised with a random number $r$, which is multiplied by the search space upper bound $\textbf{x}^{U}$ (to ensure the newly generated individual being within the search space) and divided by the current iteration step $i$ (to progressively reduce the extent of the expansion in the search space and guaranteeing convergence). Note that the iteration step can also be multiplied by a further parameter to adjust the explosion strength and quicken convergence [45].
$\textbf{x}^{new}=c+\frac{\textbf{x}^{U}r}{i},\quad\textbf{x}\in P$ (4)
II-B Clustering for Multi-Modal Optimization
The metaheuristic of reducing individuals to their center of mass makes BBBC the perfect candidate for being a multi-modal optimization solver based on clustering [38]. The hypothesis is that if a single run of BBBC can crunch to the global optimum, then multiple parallel runs on separate regions of the search space can crunch to every local optimum by forming clusters of solutions.
@@ II-B Clustering for Multi-Modal Optimization
  • each sub-population defined by a cluster is crunched to its center of mass (line 8) – although k-means returns the set of centers of mass evaluated during clustering, for k-BBBC we used the crunch operator expressed in (3), which selects the best-fitting individual per cluster; and
  • in the following iteration, each center of mass is expanded to generate a new population of individuals around the $k$ centers of mass (line 5) – the bang operation is performed as in (4).
@@
5                   $\textnormal{{P}}\leftarrow\textnormal{{bang}}(\textnormal{{C}},i)$;
6
[…]
10
11      return C;
12
Algorithm 1 k-BBBC
Fig. 1 shows a typical behavior of k-BBBC through three generations (early, mid, and late) on a function with four local minima. The figure does not indicate which individual belongs to which cluster, as it is irrelevant to the algorithm; however, at the end of the run, all points visibly converged to every local minimum – regardless of their cluster.
II-C Parameter Setting and Unexpected Behaviors
During development, we observed an unexpected outcome from k-means: sometimes, a well-defined set of points is erroneously separated into two clusters, resulting in two spatially-separated sets of points to be merged in a single cluster as k-means is forced to divide the dataset into exactly ‘$k$’ clusters (a graphical description is depicted in Fig. 2). This might happen because k-means cannot adapt to different cluster densities, even when the clusters are spherical, have equal radii, or are well-separated [47]. We observed that over 100 runs of k-BBBC on the function depicted in Fig. 2 (which is very simple), k-means has a $61\%$ failure rate, resulting in k-BBBC retrieving only three of the four local optima six times over ten. We also tested an alternative clustering method, k-medoids [48], which differs from k-means in terms of the points chosen to be the centroids of the clusters (in this case, namely the medoids): k-means takes the mean of the dataset, whereas k-medoids uses actual points from the dataset [49]. Although we observed a lower failure rate on the same settings ($21\%$), k-medoids has a higher computational complexity and does not solve the problem completely.
(a) Correct
(b) Failure
(c) Loss
To prevent uncontrolled outcomes produced by user-defined parameters, we offer a formalization for the relationship between $k$ and $m$. We found that the right balance is to account for two clusters per dimension for each local optimum of the problem, setting the value of $k$ as defined in (5). As shown in Fig. 3, this relationship provides the right balance to obtain centers of masses crunched in the very proximity of their respective local optimum. By setting $k<m$, some optima can be missed (due to the effect reported in Fig. 2); on the other hand, setting $k\gg m$ will cause the accumulation of centers of mass in a single concave region due to overpopulation, deceiving the post-processing methods described in Sec. III.
$k=2\cdot m\cdot d$ (5)
Lastly, we defined a method for determining the population size $n$ depending on the problem to be optimized. In an EA, the population size plays a pivotal role in determining its ability to explore the search space (diversity) and the time it takes to find a satisfactory solution (convergence speed). Adjusting this value is crucial to achieve a balance between solution quality and computational efficiency. Since k-BBBC works by dividing the population into different clusters, the population size must be a multiple of the number of clusters $k$; if the population size is arbitrarily chosen by the user, there is the risk that clusters might not be sufficiently populated. We empirically estimated a multiplication factor for determining $n$ from $k$, starting from small values (i.e., $5$) and progressively increasing it. We found that a value of $20$ was sufficient to explore the search spaces of the tested functions and to correctly converge to local optima. Therefore, the expression in (6) allows the population size to vary based on the problem to be optimized, as it depends on the number of clusters and the number of optima. Higher values than $20$ might also be used to increase exploration to the detriment of runtime.
$n=20\cdot k$ (6)
(a) $k=m$
(b) $k=5m$
(c) $k=2m$
Figure 3: When setting the value of $k$, it is suggested to divide each concave upward/downward region among at most two clusters per dimension of the problem. When $k$ is underestimated, some optima can be lost; when overestimated, the post-processing methods can be deceived (Sec III). In this example, we ran k-BBBC on a problem with $m=24$ local minima, $g=1000$ generations, and a population of $n=20\cdot k$ individuals. In (a), an example of underestimation with $k=m=24$ producing errors similar to the one reported in Fig. 2, resulting in missing optima. In (b), an example of overestimation with $k=5\cdot m=120$: this results in the population not being completely flattened at the end of the run, and since the population distribution is not uniform through iterations, some clusters will necessarily share part of the search space. In (c), a correct estimation with $k=2\cdot m=48$. The centers of mass correctly flattened at the local minima, with repetitions barely visible due to the close proximities of centers of mass within the same concave upward region.
II-D Elitism
output : Array of centers of mass C.
1 begin
[…]
3       for $i\in[1,\textnormal{{g}}]$ do
4             if $i\neq 1$ then
5                   $\textnormal{{P}}\leftarrow\textnormal{{bang}}(\textnormal{{C}},i)$;
[…]

### v2

# Multimodal Optimization with k-Cluster Big Bang-Big Crunch Algorithm and Postprocessing Methods for Identification and Quantification of Optima

@@ Abstract
Multimodal optimization is often encountered in engineering problems, especially when different and alternative solutions are sought. Evolutionary algorithms can efficiently tackle multimodal optimization thanks to their features such as the concept of population, exploration/exploitation, and being suitable for parallel computation. This paper investigates whether a less-known optimizer, the Big Bang-Big Crunch (BBBC) algorithm, is suitable for multimodal optimization. We extended BBBC and propose k-BBBC, a clustering-based multi-modal optimizer. Additionally, we introduce two post-processing methods to (i) identify the local optima in a set of retrieved solutions (i.e., a population), and (ii) quantify the number of correctly retrieved optima against the expected ones (i.e., success rate). Our results show that k-BBBC performs well even with problems having a large number of optima (tested on $379$ optima) and high dimensionality (tested on $32$ decision variables), but it becomes computationally too expensive for problems with many local optima (i.e., in the CEC’2013 benchmark set). Compared to other multimodal optimization methods, it outperforms them in terms of accuracy (in both search and objective space) and success rate (number of correctly retrieved optima) when tested on basic multimodal functions, especially when elitism is applied; however, it requires knowing the number of optima of a problem, which makes its performance decrease when tested on niching competition test CEC’2013. Lastly, we validated our proposed post-processing methods by comparing their success rate to the actual one: results suggest that these methods can be used to evaluate the performance of a multimodal optimization algorithm by correctly identifying optima and providing an indication of success – without the need to know where the optima are located in the search space.
keywords:
Multimodal Optimization , Evolutionary Algorithms , Clustering, Big Bang-Big Crunch
PACS:
87.85.St , 87.55.de , 87.55.kd
MSC:
49
journal: ArXiv Preprint
[…]
[inst1]organization=Computer Engineering Department, Kadir Has University, addressline=Cibali, Kadir Has Cd., Fatih, city=Istanbul, postcode=34083, country=Turkey
1 Introduction
@@
Mathematical optimization is the search for the best element within a set of alternatives based on specific criteria, with applications that range in many different areas such as science, economics, and engineering [1, 2, 3, 4, 5, 6]. The process usually consists of exploring the search space of a given objective function, which could be limited by well-defined constraints, to retrieve a single global optimal solution that either minimizes or maximizes the objective – namely, global optimization. When the objective function features more than one suitable solution (i.e., optima or modes), the process is known as multimodal optimization.
Many practical optimization problems have more than one optimal solution – whether that is multiple optima with the same objective value or a single global optimum surrounded by many local optima. In such cases, retrieving a single optimum might not be sufficient to tackle the problem due to different reasons: (i) a global optimum might be defined by parameters that are unrealistic or excessively expensive [7]; (ii) a solution can be optimal based on the availability of essential resources or the fulfillment of specific criteria, but if these conditions change over time, users need to adapt to different solutions [8]; (iii) knowing various ideal solutions within the search space can offer a valuable understanding of the characteristics of the problem and its solutions [9]; (iv) robust solutions are preferred over global ones because if the optimal solution lies on a peak of the function, any small change in its value would cause instability to the system, whereas a sub-optimal solution lying in a wider area or a plateau is less susceptible to change [10]; and (v) reliable solutions are preferred over global ones, as uncertainty in the search space might lead to infeasible optimal solutions (i.e., in violation of constraints) [11]. Under these circumstances, a local optimum offering decent performance at a reasonable cost might be more desirable than a costly global optimum providing a marginal better performance [12].
@@
Evolutionary Computation offers a set of optimization algorithms, namely Evolutionary Algorithms (EAs), which can easily tackle complex optimization problems such as multimodal ones. Thanks to their stochastic exploration of space, along with the inherent parallel computation offered by their population nature, they can easily locate different good solutions at the same time [13, 14]. Many multimodal evolutionary algorithms (MMEAs) have been proposed in the literature, featuring diverse strategies such as: (i) implicitly promote diversity by spatial segregation [15, 16] or spatial distribution [17, 18, 19, 20, 21, 22] of a single population; (ii) partitioning the population [23] and imposing mating restrictions between different populations [24, 25] or depending on the individuals’ life cycle [26]; (iii) with elitist methods [27] or methods that conserve the genotype of individuals [28]; (iv) enforcing diversity with fitness sharing [29, 30, 31], or clearing [32], crowding [33, 34, 35, 36, 37, 38], clustering [39]; and (v) employing multi-objective optimization techniques [8]. However, most of these methods suffer from difficulties such as tuning niching parameters, poor scalability with high dimensionality, extra computational overhead, not being able to fully explore the search space, relying on gradient information, premature convergence, and failure to find the right balance between exploration (finding new solutions in the search space) and exploitation (making use of the features of the good solutions discovered throughout the run) [40, 41].
@@
In this work, we investigate how the efficient and yet relatively unknown Big Bang-Big Crunch (BBBC) algorithm [42] can be extended to solve multimodal optimization problems. This investigation is motivated by the advantages of BBBC over genetic algorithms: premature convergence, convergence speed, and execution time. Our proposed algorithm, namely the k-Cluster Big Bang-Big Crunch algorithm (k-BBBC), is a method that relies on uniform exploration of the search space, as well as on exploitation of its best solutions – with the downside of knowing an estimation of the number of optima sought. An additional strength of BBBC is to dynamically turn exploration into exploitation during a single run, guaranteeing convergence to an optimal set of solutions; we used this property, allowing k-BBBC to correctly and fully converge to the local optima of a problem – note that the entire population will converge. However, the amount of initial exploration must be sufficient to find all the optima [43, 44]. To ensure a proper amount of exploration, we mathematically defined the relationship between the population size and the number of local optima to be retrieved (when known), which allowed k-BBBC to find all optima for every problem we tested – with some minor exceptions for basic multimodal functions, and for large-scale problems that require more advanced techniques.
@@
However, returning a population of converged solutions without identifying which ones are the optima is not efficient (i.e., we cannot simply sort solutions based on their fitness). We certainly do not want to manually compare solutions against each other: this quickly becomes unfeasible when the population size and the number of expected optima increases. Therefore, we propose efficient and automated methods for post-processing the output of an MMEA – theoretically, every MMEA provided that its population fully converged. As additional contributions, we propose the following two clustering-based methods:
  • 1.
    an identification procedure to detect the local optima in a population of converged solutions, and
  • 2.
    a quantification procedure to determine how many local optima have been missed (when this information is available).
The rest of the paper is organized as follows. Sec. 2 describes k-BBBC and Sec. 3 describes post-processing procedures for precise localization of optimal solutions within a wide population of individuals. Then, we performed two experiments. The first, in Sec. 4, tests k-BBBC on basic multimodal functions, with dimensionality ranging from 1 to 32, showing that the algorithm correctly converges (in most cases) to all local optima; in this experiment, k-BBBC was compared with the only publicly available MMEA on Mathworks File Exchange, MCS [7], outperforming it. In this section, we also prove that the identification and quantification methods work during post-processing by comparing their results with the known solutions. The second experiment, in Sec. 5, compares k-BBBC to thirteen state-of-the-art MMEAs developed in the last decade on the CEC’2013 benchmark set [45]: this benchmark, however, is designed for algorithms that retrieve multiple global optima only and ignore any local optima – which are considered as traps. Its functions mostly feature a large and unknown number of local minima, which cannot be estimated, making k-BBBC unsuitable for it – i.e., k-BBBC needs an estimation of the number of optima sought. Nevertheless, the results are surprisingly acceptable when compared with the latest state-of-the-art, considering that the algorithm was not designed for this purpose. Sec. 6 discusses the advantages and disadvantages of k-BBBC based on the experimental results, leaving room for several future works and opportunities that will be discussed in Sec. 7.
2 Optimization Method
@@
This section will present our proposed multimodal optimization method, starting from its original version for global optimization to a full walk-through of our proposed method and its behavior. We will be using the function defined in Eqn. (1), namely the Key function, to provide visual reference while describing the algorithm. This function features a single decision variable $x$ and a changeable number of optima depending on the parameter $m$ – featuring only one global minimum. Lastly, Table 1 reports the list of the algorithm’s parameters along with the symbol we will use to refer to them in the following sections.
@@
Table 1: Algorithm’s Parameters
@@
$g$ max number of iterations
$m$ number of local optima
$k$ number of clusters
@@
$d$ problem dimensionality
2.1 Original Big Bang-Big Crunch Algorithm
@@
BBBC features different crunch operators [46], which determine how the center of mass $c$ is calculated. It can be done by weighting the individuals with corresponding fitness evaluations or by choosing the most fitting individual as in Eqn. (2).
$c=\begin{dcases}x_{j}:f(x_{j})=\max_{i=1}^{n}{f(x_{i})}\;\text{if maximization}\\ x_{j}:f(x_{j})=\min_{i=1}^{n}{f(x_{i})}\;\text{if minimization}\end{dcases},\;x_{i}\in P$ (2)
The state-of-the-art reports only one type of bang operator [42], which generates new individuals $x^{new}$ for the next population as expressed in Eqn. (3). The center of mass $c$ is noised with a random number $r$, which is multiplied by the search space bounds $x^{U}$ and $x^{L}$ (to ensure the newly generated individual being within the search space) and divided by the current iteration step $i$ (to progressively reduce the extent of the expansion in the search space and guaranteeing convergence). Note that the iteration step can also be multiplied by a further parameter to adjust the explosion strength and quicken convergence [47].
$x^{new}=c+\frac{(x^{U}-x^{L})\cdot r}{i},\quad x\in P$ (3)
2.2 Clustering for Multimodal Optimization
The metaheuristic of reducing individuals to their center of mass makes BBBC the perfect candidate for being a multimodal optimization solver based on clustering [39]. The hypothesis is that if a single run of BBBC can crunch to the global optimum, then multiple parallel runs on separate regions of the search space can crunch to every local optimum by forming clusters of solutions.
@@ 2.2 Clustering for Multimodal Optimization
  • 3.
    each sub-population defined by a cluster is crunched to its center of mass (line 8) – although k-means returns the set of centers of mass evaluated during clustering, for k-BBBC we used the crunch operator expressed in Eqn. (2), which selects the best-fitting individual per cluster; and
  • 4.
    in the following iteration, each center of mass is expanded to generate a new population of individuals around the $k$ centers of mass (line 5) – the bang operation is performed as in Eqn. (3).
@@
6                   $\textnormal{{P}}\leftarrow\textnormal{{bigBang}}(\textnormal{{C}},i)$;
@@ 8            $\textnormal{{P}}\leftarrow\textnormal{{evaluation}}(f,\textnormal{{P}})$;
Algorithm 1 k-BBBC
Figure 1 shows a typical behavior of k-BBBC through three generations (early, mid, and late) on a function with four local minima. The figure does not indicate which individual belongs to which cluster, as it is irrelevant to the algorithm; however, at the end of the run, all points visibly converged to every local minimum – regardless of their cluster.
2.3 Parameter Setting and Unexpected Behaviors
During development, we observed an unexpected outcome from k-means: sometimes, a well-defined set of points is erroneously separated into two clusters, resulting in two spatially-separated sets of points to be merged in a single cluster as k-means is forced to divide the dataset into exactly ‘$k$’ clusters (a graphical description is depicted in Figure 2). This might happen because k-means cannot adapt to different cluster densities, even when the clusters are spherical, have equal radii, or are well-separated [49]. We observed that over 100 runs of k-BBBC on the function depicted in Figure 2 (which is very simple), k-means has a $61\%$ failure rate, resulting in k-BBBC retrieving only three of the four local optima six times over ten. We also tested an alternative clustering method, k-medoids [50], which differs from k-means in terms of the points chosen to be the centroids of the clusters (in this case, namely the medoids): k-means takes the mean of the dataset, whereas k-medoids uses actual points from the dataset [51]. Although we observed a lower failure rate on the same settings ($21\%$), k-medoids has a higher computational complexity and does not solve the problem completely.
(a) Correct
[…]
To prevent uncontrolled outcomes produced by user-defined parameters, we offer a formalization for the relationship between $k$ and $m$. We found that the right balance is to account for two clusters per dimension for each local optimum of the problem, setting the value of $k$ as defined in Eqn. (4). As shown in Figure 3, this relationship provides the right balance to obtain centers of masses crunched in the very proximity of their respective local optimum. By setting $k<m$, some optima can be missed (due to the effect reported in Figure 2); on the other hand, setting $k\gg m$ will cause the accumulation of centers of mass in a single concave region due to overpopulation, deceiving the post-processing methods described in Sec. 3.
$k=2\cdot m\cdot d$ (4)
Lastly, we defined a method for determining the population size $n$ depending on the problem to be optimized. In an EA, the population size plays a pivotal role in determining its ability to explore the search space (diversity) and the time it takes to find a satisfactory solution (convergence speed). Adjusting this value is crucial to achieve a balance between solution quality and computational efficiency. Since k-BBBC works by dividing the population into different clusters, the population size must be a multiple of the number of clusters $k$; if the population size is arbitrarily chosen by the user, there is the risk that clusters might not be sufficiently populated. We empirically estimated a multiplication factor for determining $n$ from $k$, starting from small values (i.e., $5$) and progressively increasing it. We found that a value of $20$ was sufficient to explore the search spaces of the tested functions and to correctly converge to local optima. Therefore, the expression in Eqn. (5) allows the population size to vary based on the problem to be optimized, as it depends on the number of clusters and the number of optima. Higher values than $20$ might also be used to increase exploration to the detriment of runtime.
$n=20\cdot k$ (5)
(a) %****␣manuscript.tex␣Line␣425␣****$k=m$
[…]
Figure 3: When setting the value of $k$, it is suggested to divide each concave upward/downward region among at most two clusters per dimension of the problem. When $k$ is underestimated, some optima can be lost; when overestimated, the post-processing methods can be deceived (Sec 3). In this example, we ran k-BBBC on a problem with $m=24$ local minima, $g=1000$ generations, and a population of $n=20\cdot k$ individuals. In (a), an example of underestimation with $k=m=24$ producing errors similar to the one reported in Figure 2, resulting in missing optima. In (b), an example of overestimation with $k=5\cdot m=120$: this results in the population not being completely flattened at the end of the run, and since the population distribution is not uniform through iterations, some clusters will necessarily share part of the search space. In (c), a correct estimation with $k=2\cdot m=48$. The centers of mass correctly flattened at the local minima, with repetitions barely visible due to the close proximities of centers of mass within the same concave upward region.
2.4 Elitism
output : Array of centers of mass C.
1
2begin
[…]
4       for $i\in[1,\textnormal{{g}}]$ do
5             if $i\neq 1$ then
6                   $\textnormal{{P}}\leftarrow\textnormal{{bigBang}}(\textnormal{{C}},i)$;
[…]
8
[…]
12
13      return C;
14